A postal clerk sold 75 stamps, some of which were 5 cents and some were 8 cents. The total amount was $4.80. How many of each stamp was sold?
Answer by checkley75(3666)

E+F=75 E=75-F
8E+5F=480
8(75-F)+5F=480
600-8F+5F=480
-8F+5F=480-600
-3F=-120
F=-120/-3
F=40 NUMBER OF FIVE CENT STAMPS.
E+40=75
E=75-40
E=35 EIGHT CENT STAMPS
PROOF
8*35+5*40=480
280+200=480
480=480